∫{1+tanx tan(x+theta)} dx , plz as soon as possible, thank you.

Grade:12

1 Answers

Arun
6 years ago
We know that, tan(c-b) = (tan c – tan b)/(1+tan c * tan b),
So, (1+tan c * tan b) = (tan c – tan b) / tan (c-b),
Put c= x+\large \theta, b=x,
(1+tanx * tan(x+\large \theta)) = (tan(x+\large \theta) – tan x) / tan(x+\large \theta-x) = (tan(x+\large \theta) – tan x) / tan \large \theta,
Now , seperate the two terms in 2 diff integrals and solve.